# On the Low-Temperature MCMC threshold:

### v1

@@ Abstract.
Over the last years, there has been a significant amount of work studying the power of specific classes of computationally efficient estimators for multiple statistical parametric estimation tasks, including the estimators classes of low-degree polynomials, spectral methods, and others. Despite that, our understanding of the important class of MCMC methods remains quite poorly understood. For instance, for many models of interest, the performance of even zero-temperature (greedy-like) MCMC methods that simply maximize the posterior remains elusive.
@@
Interestingly, in the sparse tensor PCA model our results indicate that low-temperature local MCMC methods significantly underperform compared to other studied time-efficient methods, such as the class of low-degree polynomials. Moreover, the “amount” by which they underperform appears connected to the size of the computational-statistical gap of the model. Specifically, an intriguing mathematical pattern is observed in sparse tensor PCA, which we call “the geometric rule”: the information-theory threshold, the threshold for low-degree polynomials and the MCMC threshold form a geometric progression for all values of the sparsity level and tensor power.
@@ 1.1.2. Sparse Tensor PCA
$\lambda_{\mathrm{STATS}}(t,k)=\widetilde{\Theta}\left(\sqrt{k}\right)$
where the $\widetilde{\Theta}$ notation here and throughout the paper hides $\mathrm{poly}\log p$-dependent terms.
@@
$\lambda_{\mathrm{ALG}}(t,k)=\widetilde{\Theta}\left(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\}\right).$
@@
Yet, a quick check in the proof of their lower bound [LZ22, Theorem 16] shows that in fact the authors first prove the conditional failure of polynomial-time algorithms when $\lambda=o(\lambda_{\mathrm{ALG}}(t,k))$ for our case of interest, that is when $v_{1}=\ldots=v_{t}$, and then add an extra step to also prove it for the case that $v_{i},i=1,\ldots,t$ are independent. In particular, their result applies almost identically to the case of symmetric sparse tensor PCA we focus on this work. For reasons of completeness, in Section 11 we describe exactly the slightly modified argument from [LZ22] and how it applies in our symmetric setting.
@@
$\lambda_{\mathrm{MCMC}}(t,k)=\widetilde{\Theta}\left(\min\left\{k^{t-\frac{1}{2}},\,\frac{p^{t-1}}{k^{t-\frac{3}{2}}}\right\}\right).$
@@
Satisfyingly, we also prove that when $\lambda=\widetilde{o}(\lambda_{\mathrm{MCMC}}(t,k))$ a negative result for all low temperature reversible local MCMC methods (including the Metropolis chain) holds. Specifically, we prove in Theorem 3.6 that for large enough $\beta>0$ all local reversible Markov chains with stationary measure $\pi_{\beta}$ mix in super-polynomial time. Notably, our result significantly improves upon the prior work [AWZ23] on sparse PCA where a similar negative result for local Markov chains was proven but only when $t=2$ and importantly under the more restrictive condition $\lambda=o(\lambda_{\mathrm{ALG}}(2,k)).$ The proof of Theorem 3.6 is the most technical contribution of the present work. We prove – via a delicate conditional second moment method – a bottleneck for the MCMC dynamics based on the Overlap Gap Property framework from [GZ22a]. The proof carefully leverages the “flatness” idea from random graph theory [BBSV19, GZ19], but adjusted to high dimensional Gaussian tensors, and could be of independent interest.
@@
$\sqrt{\lambda_{\mathrm{MCMC}}(t,k)\times\lambda_{\mathrm{STATS}}(t,k)}=\widetilde{\Theta}(\lambda_{\mathrm{ALG}}(t,k)).$ (1.3)
@@ 1.2. Asymptotic Notation
We will use the standard asymptotic notations $O,o,\Omega,\omega,\Theta$. On top of that, for any $(a_{p})_{p\in\mathbb{N}}$, $(b_{p})_{p\in\mathbb{N}}$ sequences of real numbers, we write $a_{p}=\widetilde{o}(b_{p})$ when $a_{p}=o(b_{p}(\log p)^{x})$ for some $x\in\mathbb{Z}$, $a_{p}=\widetilde{\omega}(b_{p})$ if $b_{p}=\widetilde{o}(a_{p})$ and $b_{p}=\widetilde{\Theta}(a_{p})$ if $\Omega(b_{p}(\log p)^{y})=a_{p}=O(b_{p}(\log p)^{x})$ for some $y,x\in\mathbb{Z}$.
@@ 3.3. The negative result
Theorem 3.6 is proven in Section 9.
@@ 4.3. The negative result
$\beta>\frac{2\log(p/k)}{n}\exp\left(\frac{2k\log(p/k)}{n}\right)$
@@ 5.2.1. Specification
$n_{STATS}=\frac{2k\log p}{\log\left(\frac{k}{\sigma^{2}}+1\right)}\approx 61$
@@
$n_{LASSO}=(2k+\sigma^{2})\log p\approx 357$
@@ 5.2.2. Results
The results of these simulations are displayed in Fig. 4. We again observe a clear phase transition when the sample size $n$ crosses the algorithmic threshold. When $n<n_{LASSO}$, most samples time out after a million iterations having failed to find $v^{*}$. Above $n_{LASSO}$, the chain quickly finds $v^{*}$. We can see these dynamics in more detail in Fig. 5. Here we group simulations by their value of $n$. Then at each time, we plot the median proportion of $v^{*}$ recovered at that time within that group. We observe that for values of $n$ above 300, we achieve very rapid convergence. Conversely, below 175 the chain does not appear to be converging at all, despite being significantly above the information-theoretic threshold $n_{STATS}\approx 61$.
@@
Figure 5. Sparse Regression: Median proportion of $v^{*}$ recovered over time for different values of $n$. Here $p=20000$, $k=18$, $\sigma^{2}=0.05$ and $\beta=100$. The information theoretic threshold is $\approx 61$, and the predicted algorithmic threshold is $\approx 357$.
@@ 6.2. Comparison with [GZ22a]
In our work, we prove more general both positive and negative results. For the positive side, we prove that for all $k=o(p),$ a larger family of low temperature MCMC methods succeeds whenever $n=\omega(n_{\mathrm{ALG}})$. In particular, our positive results include the randomized greedy algorithm maximizing the posterior, as opposed to the vanilla greedy approach of [GZ22a]. We highlight that this is a non-trivial extension, which is not always known to be true (see e.g., the relevant discussion in [GJX23] for the planted clique model) Finally, we prove as our negative result that for all $k=o(p^{1/4})$ a large family of low temperature MCMC methods fails when $n=o(n_{\mathrm{ALG}})$, significantly again generalizing the result of [GZ22a].
@@ 8.1. Proof of Positive Results for the Gaussian Additive Model
Here we apply Theorem 2.3 to prove Theorem 3.3 under 3.2. The first step is to show that 3.2 implies 2.2.
8.1.1. Verifying 2.2
@@
In the following lemma, we verify 2.2 under 3.2.
@@
Assume that 3.2 holds with constants $\varepsilon\in(0,1)$ and $t\geq 1$. For all positive integers $p,k,\ell_{\min}\in\mathbb{N}^{+}$ such that $2/\varepsilon\leq k\leq p$ and $\ell_{\min}\leq\varepsilon k$, if
@@
Fix $x\in\mathcal{X}_{p,k}$, and suppose $|x\cap v^{*}|=\ell$ for some integer $\ell\leq k-1$. Hence, we have $|y\cap v^{*}|=\ell+1$ for all $y\in\Gamma_{G,v^{*}}(x)$. Observe from Eq. 3.2 that
@@
We deduce from 3.2 that
@@
where we use the convexity of $f$ and item (b) of 3.2. Thus, we have
@@
where we apply the convexity of $f$ and item (c) of 3.2. Meanwhile, the number of $y^{\prime}\in\Gamma_{G,v^{*}}(x)$ such that $\left\langle y,y^{\prime}\right\rangle=k-1$ is $2(k-\ell-1)$, and for such $w$ it holds
@@
by the convexity of $f$ and item (b) of 3.2. Finally, if $y^{\prime}=y$ then $\left\langle y,y^{\prime}\right\rangle=k$ and
@@
We are now ready to prove Theorem 3.3. We first consider the sparse case where $k=O(\sqrt{p\log p})$, where we directly apply Theorems 2.3 and 8.1. We need to show that 2.2 is satisfied with high probability, and then the results follow from Theorem 2.3. By Lemma 8.1 with $\ell_{\min}=0$, for every $x\in\mathcal{X}_{p,k}$ we pick an arbitrary $y\in\Gamma_{G,v^{*}}(x)$ such that
@@
and set the parent of $x$ to be $y$; this gives a spanning tree with high probability. Note that this automatically establishes Eq. 2.1 from 2.2 with $\delta=\Omega(\lambda/k^{t})$. Observe that the diameter of the tree is at most $2k$, and the maximum degree of the graph $G_{p,k}$ is at most $kp$. Hence, 2.2 follows. To upper bound the running time, notice that
@@
Also recalling Eq. 3.2, the range of $H$ is at most
@@
Under the assumptions of Theorem 3.3 and suppose $k=\omega(\sqrt{p\log p})$, if we initialize the Metropolis chain with a uniformly random $X_{0}\in\mathcal{X}_{p,k}$, then with high probability it holds $|X_{t}\cap v^{*}|\geq\frac{k^{2}}{10p}$ for all $t\leq e^{O(k^{2}/p)}$.
@@
for Eq. 2.1 in 2.2. Thus, by Theorem 2.3, for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-2}}\log p)$ the number of iterations required is at most $O(\beta\lambda k^{2}p)$ with high probability, and for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-3}}\log(p/k))$ it is at most
@@ 8.2. Proof of Positive Results for Sparse Linear Regression
Here we prove Theorem 4.2 under 4.1. First we show that 4.1 implies 2.2.
8.2.1. Verifying 2.2
In order to verify 2.2 we present a Binary - Local Search Algorithm that is a binary version of the Local Search Algorithm in [GZ22a]. The proof of its convergence is essentially the same as for the Local Search Algorithm, except we only need to consider support recovery. We use $e_{1},e_{2},\dots$ to refer to the standard basis vectors in $\mathbb{R}^{p}$. We write $X_{i}$ to refer to the $i^{th}$ column of $X$.
@@
B-LSA looks at all the possible coordinate “flips” we could make in one step, and chooses the one that minimises the squared error. It repeats until no local improvement is possible. In the following lemma, we verify the second part of 2.2 under 4.1, with $\delta=\frac{n}{4}$ by showing that B-LSA always reduces the squared error by a factor of at least order $n$ at each step.
@@
Assume that 4.1 holds. If $v\neq v^{*}$, then for any $v^{\prime}$ obtained from $v$ in one iteration of B-LSA, with high probability it holds that
@@
The proof of Lemma 8.4 is deferred until Section 12. In the next lemma, we verify the first part of 2.2 with $D=O(k)$.
@@
Assume that 4.1 holds. With high probability as $k\rightarrow\infty$, B-LSA with uniform random initialisation $v_{0}$ terminates in at most $O(k)$ iterations and outputs $v^{*}$.
@@
Note that since $\sigma^{2}\leq ck$ by 4.1, we have that $20\left(2(k-\ell)+\sigma^{2}\right)\leq(40+20c)k$, and therefore
@@
We assume 4.1 and $n\geq Ck\log(p/k)$ for some $C>0.$ Then with high probability as $k\rightarrow\infty$, it holds that
@@
Therefore 2.2 is satisfied with $D=O(k)$ and $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn).$ We now have everything we need to prove Theorem 4.2.
@@
We note that in this case that $\Delta$, the maximum degree of the graph, is given by $\Delta=\max_{v\in\mathcal{X}_{p,k}}|\Gamma_{G}(v)|=k(p-k)$, as each $v$ has $k$ 1’s and $(p-k)$ 0’s. By Lemma 8.4, Lemma 8.5, and Lemma 8.6, we have that 2.2 holds with $D=O(k)$, $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn)$. We also see that the cardinality of the state space $\mathcal{X}$ is given by $\binom{p}{k}$. We therefore apply the first two parts of Theorem 2.3 to see that if

### v2

@@ Abstract.
Over the last years, there has been a significant amount of work studying the power of specific classes of computationally efficient estimators for multiple statistical parametric estimation tasks, including the estimator classes of low-degree polynomials, spectral methods, and others. Despite that, our understanding of the important class of MCMC methods remains quite poorly understood. For instance, for many models of interest, the performance of even zero-temperature (greedy-like) MCMC methods that simply maximize the posterior remains elusive.
@@
Notably, in the sparse tensor PCA model our results indicate that low-temperature local MCMC methods significantly underperform compared to other studied time-efficient methods, such as the class of low-degree polynomials. Moreover, the “amount” by which they underperform appears connected to the size of the computational-statistical gap of the model. Specifically, an intriguing mathematical pattern is observed in sparse tensor PCA, which we call “the geometric rule”: the information-theory threshold, the threshold for low-degree polynomials and the MCMC threshold form a geometric progression for all values of the sparsity level and tensor power.
@@ 1.1.2. Sparse Tensor PCA
$\lambda_{\mathrm{STATS}}(t,k)=\tilde{\Theta}\left(\sqrt{k}\right)$
where the $\tilde{\Theta}$ notation here and throughout the paper hides $\mathrm{poly}\log p$-dependent terms.
@@
$\lambda_{\mathrm{ALG}}(t,k)=\tilde{\Theta}\left(\min\{k^{t/2},p^{(t-1)/2}/k^{t/2-1}\}\right).$
@@
Yet, a quick check in the proof of their lower bound [LZ22, Theorem 16] shows that in fact the authors first prove the conditional failure of polynomial-time algorithms when $\lambda=o(\lambda_{\mathrm{ALG}}(t,k))$ for our case of interest, that is when $v_{1}=\ldots=v_{t}$, and then add an extra step to also prove it for the case that $v_{i},i=1,\ldots,t$ are independent. In particular, their result applies almost identically to the case of symmetric sparse tensor PCA we focus on this work. For reasons of completeness, in Section 11, we describe exactly the slightly modified argument from [LZ22] and how it applies in our symmetric setting.
@@
$\lambda_{\mathrm{MCMC}}(t,k)=\tilde{\Theta}\left(\min\left\{k^{t-\frac{1}{2}},\,\frac{p^{t-1}}{k^{t-\frac{3}{2}}}\right\}\right).$
@@
Satisfyingly, we also prove that when $\lambda=\tilde{o}(\lambda_{\mathrm{MCMC}}(t,k))$ a negative result for all low temperature reversible local MCMC methods (including the Metropolis chain) holds. Specifically, we prove in Theorem 3.6 that for large enough $\beta>0$ all local reversible Markov chains with stationary measure $\pi_{\beta}$ mix in super-polynomial time. Notably, our result significantly improves upon the prior work [AWZ23] on sparse PCA where a similar negative result for local Markov chains was proven but only when $t=2$ and importantly under the more restrictive condition $\lambda=o(\lambda_{\mathrm{ALG}}(2,k)).$ The proof of Theorem 3.6 is the most technical contribution of the present work. We prove – via a delicate conditional second moment method – a bottleneck for the MCMC dynamics based on the Overlap Gap Property framework from [GZ22a]. The proof carefully leverages the “flatness” idea from random graph theory [BBSV19, GZ19], but adjusted to high dimensional Gaussian tensors, and could be of independent interest.
@@
$\sqrt{\lambda_{\mathrm{MCMC}}(t,k)\times\lambda_{\mathrm{STATS}}(t,k)}=\tilde{\Theta}(\lambda_{\mathrm{ALG}}(t,k)).$ (1.3)
@@ 1.2. Asymptotic Notation
We will use the standard asymptotic notations $O,o,\Omega,\omega,\Theta$. On top of that, for any $(a_{p})_{p\in\mathbb{N}}$, $(b_{p})_{p\in\mathbb{N}}$ sequences of real numbers, we write $a_{p}=\tilde{o}(b_{p})$ when $a_{p}=o(b_{p}(\log p)^{x})$ for some $x\in\mathbb{Z}$, $a_{p}=\tilde{\omega}(b_{p})$ if $b_{p}=\tilde{o}(a_{p})$ and $b_{p}=\tilde{\Theta}(a_{p})$ if $\Omega(b_{p}(\log p)^{y})=a_{p}=O(b_{p}(\log p)^{x})$ for some $y,x\in\mathbb{Z}$.
@@ 3.3. The negative result
Theorem 3.6 is proved in Section 9.
@@ 4.3. The negative result
$\beta>\Omega\left(\frac{\log(p/k)}{n}\exp\left(\frac{2k\log(p/k)}{n}\right)\right)$
@@ 5.2.1. Specification
$n_{STATS}=\frac{2k\log(p/k)}{\log\left(\frac{k}{\sigma^{2}}+1\right)}\approx 43$
@@
$n_{LASSO}=(2k+\sigma^{2})\log(p/k)\approx 253$
@@ 5.2.2. Results
The results of these simulations are displayed in Fig. 4. We again observe a clear phase transition when the sample size $n$ crosses the algorithmic threshold. When $n<n_{LASSO}$, most samples time out after a million iterations having failed to find $v^{*}$. Above $n_{LASSO}$, the chain quickly finds $v^{*}$. We can see these dynamics in more detail in Fig. 5. Here we group simulations by their value of $n$. Then at each time, we plot the median proportion of $v^{*}$ recovered at that time within that group. We observe that for values of $n$ above 300, we achieve very rapid convergence. Conversely, below 175 the chain does not appear to be converging at all, despite being significantly above the information-theoretic threshold $n_{STATS}\approx 43$.
@@
Figure 5. Sparse Regression: Median proportion of $v^{*}$ recovered over time for different values of $n$. Here $p=20000$, $k=18$, $\sigma^{2}=0.05$ and $\beta=100$. The information theoretic threshold is $\approx 43$, and the predicted algorithmic threshold is $\approx 253$.
@@ 6.2. Comparison with [GZ22a]
In our work, we prove more general both positive and negative results. For the positive side, we prove that for all $k=o(p),$ a larger family of low temperature MCMC methods succeeds whenever $n=\omega(n_{\mathrm{ALG}})$. In particular, our positive results include the randomized greedy algorithm maximizing the posterior, as opposed to the vanilla greedy approach of [GZ22a]. We highlight that this is a non-trivial extension, which is not always known to be true (see e.g., the relevant discussion in [GJX23] for the planted clique model) Finally, we prove as our negative result that for all $k=o(p^{1/3})$ a large family of low temperature MCMC methods fails when $n=o(n_{\mathrm{ALG}})$, significantly again generalizing the result of [GZ22a].
@@ 8.1. Proof of Positive Results for the Gaussian Additive Model
Here we apply Theorem 2.3 to prove Theorem 3.3 under Assumption 3.2. The first step is to show that Assumption 3.2 implies Assumption 2.2.
8.1.1. Verifying Assumption 2.2.
@@
In the following lemma, we verify Assumption 2.2 under Assumption 3.2.
@@
Assume that Assumption 3.2 holds with constants $\varepsilon\in(0,1)$ and $t\geq 1$. For all positive integers $p,k,\ell_{\min}\in\mathbb{N}^{+}$ such that $2/\varepsilon\leq k\leq p$ and $\ell_{\min}\leq\varepsilon k$, if
@@
Fix $x\in\mathcal{X}_{p,k}$, and suppose $|x\cap v^{*}|=\ell$ for some integer $\ell\leq k-1$. Hence, we have $|y\cap v^{*}|=\ell+1$ for all $y\in\Gamma_{G,v^{*}}(x)$. Observe from (3.2) (the definition of $H$), that
@@
We deduce from Assumption 3.2 that
@@
where we use the convexity of $f$ and item (b) of Assumption 3.2. Thus, we have
@@
where we apply the convexity of $f$ and item (c) of Assumption 3.2. Meanwhile, the number of $y^{\prime}\in\Gamma_{G,v^{*}}(x)$ such that $\left\langle y,y^{\prime}\right\rangle=k-1$ is $2(k-\ell-1)$, and for such $w$ it holds
@@
by the convexity of $f$ and item (b) of Assumption 3.2. Finally, if $y^{\prime}=y$ then $\left\langle y,y^{\prime}\right\rangle=k$ and
@@
We are now ready to prove Theorem 3.3. We first consider the sparse case where $k=O(\sqrt{p\log p})$, where we directly apply Theorem 2.3 and Lemma 8.1. We need to show that Assumption 2.2 is satisfied with high probability, and then the results follow from Theorem 2.3. By Lemma 8.1 with $\ell_{\min}=0$, for every $x\in\mathcal{X}_{p,k}$ we pick an arbitrary $y\in\Gamma_{G,v^{*}}(x)$ such that
@@
and set the parent of $x$ to be $y$; this gives a spanning tree with high probability. Note that this automatically establishes the second part of Assumption 2.2 with $\delta=\Omega(\lambda/k^{t})$. Observe that the diameter of the tree is at most $2k$, and the maximum degree of the graph $G_{p,k}$ is at most $kp$. Hence, Assumption 2.2 follows. To upper bound the running time, notice that
@@
Also recalling the definition of $H$, the range of $H$ is at most
@@
Under the assumptions of Theorem 3.3, and suppose $k=\omega(\sqrt{p\log p})$, if we initialize the Metropolis chain with a uniformly random $X_{0}\in\mathcal{X}_{p,k}$, then with high probability it holds $|X_{t}\cap v^{*}|\geq\frac{k^{2}}{10p}$ for all $t\leq e^{O(k^{2}/p)}$.
@@
for the second part of Assumption 2.2. Thus, by Theorem 2.3, for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-2}}\log p)$ the number of iterations required is at most $O(\beta\lambda k^{2}p)$ with high probability, and for $\beta\lambda=\Omega(\frac{p^{t-1}}{k^{t-3}}\log(p/k))$ it is at most
@@ 8.2. Proof of Positive Results for Sparse Linear Regression
Here we prove Theorem 4.2 under Assumption 4.1. First we show that Assumption 4.1 implies Assumption 2.2.
8.2.1. Verifying Assumption 2.2
In order to verify Assumption 2.2 we present a Binary - Local Search Algorithm that is a binary version of the Local Search Algorithm in [GZ22a]. The proof of its convergence is essentially the same as for the Local Search Algorithm, except we only need to consider support recovery. We use $e_{1},e_{2},\dots$ to refer to the standard basis vectors in $\mathbb{R}^{p}$. We write $X_{i}$ to refer to the $i^{th}$ column of $X$.
@@
B-LSA looks at all the possible coordinate “flips” we could make in one step, and chooses the one that minimises the squared error. It repeats until no local improvement is possible. In the following lemma, we verify the second part of Assumption 2.2 under Assumption 4.1, with $\delta=\frac{n}{4}$ by showing that B-LSA always reduces the squared error by a factor of at least order $n$ at each step.
@@
Assume that Assumption 4.1 holds. If $v\neq v^{*}$, then for any $v^{\prime}$ obtained from $v$ in one iteration of B-LSA, with high probability it holds that
@@
The proof of Lemma 8.4 is deferred until Section 12. In the next lemma, we verify the first part of Assumption 2.2 with $D=O(k)$.
@@
Assume that Assumption 4.1 holds. With high probability as $k\rightarrow\infty$, B-LSA with uniform random initialisation $v_{0}$ terminates in at most $O(k)$ iterations and outputs $v^{*}$.
@@
Note that since $\sigma^{2}\leq ck$ by Assumption 4.1, we have that $20\left(2(k-\ell)+\sigma^{2}\right)\leq(40+20c)k$, and therefore
@@
We assume Assumption 4.1 and $n\geq Ck\log(p/k)$ for some $C>0.$ Then with high probability as $k\rightarrow\infty$, it holds that
@@
Therefore Assumption 2.2 is satisfied with $D=O(k)$ and $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn).$ We now have everything we need to prove Theorem 4.2.
@@
We note that in this case that $\Delta$, the maximum degree of the graph, is given by $\Delta=\max_{v\in\mathcal{X}_{p,k}}|\Gamma_{G}(v)|=k(p-k)$, as each $v$ has $k$ 1’s and $(p-k)$ 0’s. By Lemma 8.4, Lemma 8.5, and Lemma 8.6, we have that Assumption 2.2 holds with $D=O(k)$, $\delta=\frac{n}{4}$, and $\mathcal{R}_{H}=O(kn)$. We also see that the cardinality of the state space $\mathcal{X}$ is given by $\binom{p}{k}$. We therefore apply the first two parts of Theorem 2.3 to see that if